OpenStudy (anonymous):

Find the two digit number whose tens digit is 4 less than its units digit if the original number is 2 more than 3 times the sum of the digits.

OpenStudy (radar):

Use algebra. Let A = the Tens digit Let B = the Ones digit The we have these two equations. A=B-4 given and 10A + B = 3(A+B) + 2 Rearrange the two equations to look like: A - B = -4 and 7A-2B = 2 Now solve using substitution or "elimination" method.